# Methodology, Data Curation, Formal analysis, Experimentation, Writing - Original Draft

### v1

@@ 5 Conclusion
This work presented ”XG-NID,” a novel framework that advances the state-of-the-art in ML-based NIDS by integrating flow-level and packet-level data into a heterogeneous graph structure. By leveraging a HGNN for graph-level classification, XG-NID provides real-time inference capabilities while capturing the complex relationships inherent in network traffic. The incorporation of LLMs adds a layer of explainability, enabling the generation of human-readable explanations and actionable insights, which are critical for effective cybersecurity defense mechanisms.
Furthermore, the introduction of new flow features based on temporal information enhances the contextual understanding and explainability of the model’s predictions. Our comprehensive evaluation demonstrates that XG-NID significantly outperforms existing baseline and state-of-the-art methods, achieving an impressive F1 score of 97% in multi-class classification. This underscores the efficacy of our approach in addressing the challenges of real-time intrusion detection while providing a robust and adaptive defense mechanism. Ultimately, XG-NID sets a new benchmark in the field of NIDS by combining innovative data fusion techniques with enhanced interpretability and real-time operational capabilities, offering a promising direction for future research and practical implementations in cybersecurity.
Future work will focus on extending XG-NID to multiple datasets to ensure its robustness across diverse network environments. Additionally, we aim to integrate it into an automated system capable of making autonomous decisions and providing rapid security responses, paving the way for intelligent, self-sustaining cybersecurity solutions

### v2

@@ 4.2.2 Payload-based Attacks’ Explanation
4.2.3 Explainability Evaluation via LLM-as-a-Judge
Evaluating the quality of natural language explanations generated by LLMs remains a challenging task due to the lack of standardized metrics. Following recent works in the literature (Zheng et al., 2023; Wang et al., 2023), we adopt the LLM-as-a-judge paradigm, which has demonstrated competitive alignment with human evaluations in multiple natural language generation tasks. This approach leverages the reasoning capabilities of a larger language model to assess the correctness and clarity of generated outputs.
To operationalize this framework, we reformulate the task of explanation evaluation as a zero-shot classification problem. Specifically, we utilize ChatGPT-4o (OpenAI, 2024) as an external evaluator that receives a generated explanation and is asked to infer the most likely attack class from a predefined list of candidate labels. This approach allows us to quantify the sense-making ability of the explanation: the better the explanation, the more accurately the LLM can identify the correct attack type from the textual description alone.
Fig. 9.: Accuracy of ChatGPT-4o in identifying the correct attack class from explanations generated by LLaMA-3 (8B) and DeepSeek-V2 Lite (15.7B) across three benchmark datasets. The results reflect the coherence and clarity of the explanations, with both models achieving comparable performance. Lower accuracy on the UNSW-NB15 dataset is consistent with its lower classification performance, as shown in Figure 4.1.1
For each explanation instance, ChatGPT-4o is provided with an anonymized explanation—i.e., with the true label withheld—and prompted to select the most probable attack class from a fixed list of attack classes. An example of the prompt used is shown below:
You are a cybersecurity expert and you are given an explanation generated by a model for a network intrusion detection scenario. Based solely on this explanation, identify the most likely attack type from the following list: {Attack_Classes_Names}.
Explanation: {Anonymized_Explanation}
Your task: Choose the attack that best fits the explanation. Only Provide the Attack Class name as your output
We compute the quality of the generated explanations by comparing the attack class predicted by the LLM (acting as a zero-shot classifier) against the true label. A higher classification accuracy indicates that the explanation is more coherent, informative, and well-aligned with the underlying attack behavior.
To ensure robustness and generalizability, this evaluation is conducted across all three benchmark datasets used in our study—CIC-IoT2023, CIC-IDS2017, and UNSW-NB15. Additionally, we benchmark the performance of our LLaMA-3-based explainer against that of DeepSeek-V2 Lite (Liu et al., 2024), a significantly larger model (15.7B parameters), under the same prompt structure and evaluation setup. Fig 4.2.3 presents the classification accuracy achieved by ChatGPT-4o when judging explanations produced by both models. As shown, both LLMs generate comparably effective explanations, indicating that the outputs of our generative explainer are generally well-aligned with the semantics of the attacks. However, a relatively lower accuracy is observed on the UNSW-NB15 dataset, which correlates with its reduced classification performance as reported earlier in Figure 4.1.1. This suggests that when the base classification itself is uncertain or noisy, the corresponding explanations are less distinguishable, even for a powerful judge model.
To further assess the relative quality of the explanations produced by each model, we conducted a pairwise comparison experiment in which ChatGPT-4o was asked to evaluate two anonymized explanations for the same attack instance—one generated by LLaMA-3 and the other by DeepSeek-V2 Lite—and select which explanation is better, or declare them equally good. The prompt used for this evaluation is provided below:
You are a cybersecurity expert. You are given the true attack type and two anonymized explanations generated by different models for the same network intrusion scenario. Based solely on the quality, coherence, and informativeness of the explanations, choose which explanation better describes the attack—or state if both are equally good.
Attack Type: {Attack_Label}
Explanation A: {Explanation by Llama3}
Explanation B: {Explanation by DeepSeek V2-Lite}
Your task: Select "Explanation A", "Explanation B", or "Tie".
Using this prompt, we computed the win rate of each model across all three datasets, quantifying how often each model’s explanation was preferred over the other. The results of this comparative analysis are presented in Fig 4.2.3. This evaluation provides a deeper insight into not only whether explanations are correct, but also how well-structured and meaningful they are from the perspective of an expert judge.
Fig. 10.: Pairwise win rate comparison between explanations generated by LLaMA-3 (8B) and DeepSeek-V2 Lite (15.7B) across all three datasets, as judged by ChatGPT-4o. The LLM-judge was provided with the true attack label and two anonymized explanations and asked to select the more informative one or declare a tie. DeepSeek appears to provide more concise and better explanation of attacks across all three datasets.
@@ 5 Conclusion
This work introduced ”XG-NID”, a novel framework that bridges a critical gap in NIDS. Traditional systems have either relied heavily on flow-level information, leaving them vulnerable to payload-dependent attacks, or focused on payload data, making them susceptible to non-payload-based threats. XG-NID addresses this issue by combining flow-level and packet-level data into a heterogeneous graph structure, effectively capturing the diverse nature of network traffic. By utilizing a HGNN for graph-level classification, XG-NID exceeds in performance, providing robust detection capabilities for both payload-dependent and non-payload-dependent attacks. Moreover, the introduction of explainability through LLMs ensures that actionable insights are generated in a human-readable format, making the system accessible to users without deep technical knowledge. This feature simplifies the decision-making process, reducing reliance on experts and empowering quicker responses to potential threats.
The results show that XG-NID not only surpasses existing state-of-the-art methods but also performs consistently well across different types of attacks, achieving an impressive F1 score of 97% in multi-class classification. We evaluated the framework on three different benchmark NIDS datasets, accompanied by a detailed analysis of the generated explanations and the actionable insights provided by the LLM. The model’s adaptability and real-time inference capabilities position it as a powerful tool for modern cybersecurity defense, setting a new standard for network intrusion detection systems.
In short, XG-NID not only addresses the critical limitations of current NIDS but also introduces a framework that enhances both performance and usability. By combining innovative data fusion techniques, advanced graph-based modeling, and enhanced interpretability, it offers a promising solution for strengthening cybersecurity defenses.
Future work will focus on extending the evaluation across additional datasets to ensure greater robustness in diverse network environments. Moreover, we aim to integrate XG-NID into an autonomous system capable of making independent security decisions and providing rapid threat responses. Additionally, we plan to further enhance the explainability component by incorporating industry-tailored fine-tuning or a Retrieval-Augmented Generation (RAG) system, allowing the model to generate responses aligned with security standards and organizational standard operating procedures (SOPs). These advancements will help pave the way toward intelligent, self-sustaining cybersecurity solutions.